OpenStudy (anonymous):

Solve: 8x - 2(x + 1) = 7x + 8

jimthompson5910 (jim_thompson5910):

8x-2(x+1)=7x+8 8x-2x-2=7x+8 6x-2=7x+8 6x=7x+8+2 6x-7x=8+2 -x=8+2 -x=10 x=(10)/(-1) x=-10